$SPAD/src/input f04mbf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
n := 10
b := new(1,n,0.0)$Matrix DoubleFloat;
b(1,1) := 6
for i in 2..n-1 repeat
  b(1,i) := 4
b(1,n) := 6
precon := true
shft := 0
-- The following two parameters are not used
-- by the ASPs. They may be useful if a user
-- supplies FORTRAN code directly.
lrwork:=1
liwork:=1
rtol := 0.00001
itnlim := 100
msglvl := 1
ifail := 1
z:=new(10,10,0.0)$Matrix MFLOAT;
z(1,1):=2;z(1,2):=1;z(1,10):=3
z(10,1):=3;z(10,9):=1;z(10,10):=2
for i in 2..(n-1) repeat
  z(i,i-1):=1
  z(i,i):=2
  z(i,i+1):=1
z
m :Matrix MFLOAT:= [[2,1,0,0,0,0,0,0,0,0],[1,2,1,0,0,0,0,0,0,0],_
[0,1,2,1,0,0,0,0,0,0],[0,0,1,2,1,0,0,0,0,0],[0,0,0,1,2,1,0,0,0,0],_
[0,0,0,0,1,2,1,0,0,0],[0,0,0,0,0,1,2,1,0,0],[0,0,0,0,0,0,1,2,1,0],_
[0,0,0,0,0,0,0,1,2,1],[0,0,0,0,0,0,0,0,1,2]]
result:=f04mbf(n,b,precon,shft,itnlim,msglvl,lrwork,liwork,_
               rtol,ifail,z::ASP28(APROD),(m::ASP34(MSOLVE)))

———-

References^{[1] nothing}

4